# Covering instability for the existence of positive scalar curvature metrics

### v1

@@ Abstract.
We show that a closed non-orientable $3$-manifold admits a positive scalar curvature metric if and only if its orientation double cover does; however, for each $4\leq n\leq 7$, there exist infinitely many non-homeomorphic smooth non-orientable $n$–manifolds such that each of them does not admit any Riemannian metric with positive scalar curvature, but its orientation double cover does. These examples were first introduced by Alpert–Balitskiy–Guth in the study of Urysohn widths.
To prove the nonexistence result, we extend the Schoen–Yau inductive descent approach to non-orientable manifolds. We also discuss band width estimates and the notion of enlargeability for non-orientable PSC manifolds.
@@ 1. Introduction
For $4\leq n\leq 7$, there exist infinitely many mutually non-homeomorphic smooth closed non-orientable manifolds with dimension $n$ such that each of them does not admit any Riemannian metric with positive scalar curvature, but its orientation double cover does.
Previously, there have been examples where a closed orientable manifold $X$ does not admit any PSC metric while a suitable finite cover of $X$ does. The first such example - an exotic $\mathbb{R}P^{9}$ was discovered by Bérard Bergery in [5, Example 9.1]. LeBrun [^1] Manifolds in [20] actually depict a sign change on the Yamabe invariants after taking double cover. [20] and Hanke–Kotschick–Wehrheim [15] constructed more examples of this type in four dimensions (see also a subsequent work [1]). All these examples have finite fundamental groups, and are closely related to a conjecture of Rosenberg [24, 1.2]. We emphasize here that orientability is crucial in these examples, in order for index-theoretic invariants (Hitchin’s $\alpha$-invariants in [5] and Seiberg–Witten in [20, 15]) to be applied to rule out the existence of PSC metrics. In comparison, each example in Theorem 1.1 has the feature that any manifold homeomorphic to it admits no PSC metric. It would be interesting to further compare these approaches.
@@ 1.1. A brief description of the construction
We describe a construction for $n=4$ and $6$. More generally, taking product with a torus $\mathbb{T}^{l}$ gives more of such examples, as long as the result is of dimensions $\leq 7$. The example described here was first introduced by Alpert–Balitskiy–Guth [2, Example 11], based on a classical example by Gromov [10, $H_{1}^{\prime\prime}$].
The original construction in [2] gives a metric space with very interesting metric-geometric properties (most notably Urysohn widths). By adding a generic perturbation, this metric space can be modified to a smooth manifold with similar properties. However, in order to understand differential geometric invariants - e.g. scalar curvature - we need to have a more precise control on the topology of the manifold. Therefore, instead of directly following the construction in [2], we will take a different approach by defining the desired manifold as a piecewise linear (PL) manifold. It is a classical result in geometric topology that every PL manifold with dimension $\leq 6$ has a unique smoothing (see [23, Theorem 2]). Working in the PL category allows us to define the desired manifold explicitly without taking any perturbations. This is crucial for the proof of the existence of PSC metrics on the orientation double cover.
@@
We will see in Section 2 that $X$ is non-orientable because the translation by $v_{2k+1}$ swaps $Z,Z^{\prime}$ and hence reverses the orientation of $X_{0}$. When $k=2,3$, define the smooth structure on $X$ to be the unique smoothing of $X$ as a PL manifold. We will show that the manifolds $X\times\mathbb{T}^{l}$ admit no PSC metrics as long as its dimension is $\leq 7$, but the orientation double cover admits PSC metrics.
@@ 1.2. Sketch of the proof
To show that $X\times\mathbb{T}^{l}$ does not admit any PSC metric, we observe that the translations of $v_{1},\cdots,v_{n-1}$ give rises to integral cohomology classes $\alpha_{1},\cdots,\alpha_{n-1}\in H^{1}(X;\mathbb{Z})$ satisfying
@@ 1.4. Remarks on higher dimensions
In our proof of the nonexistence result for $X$, we work in low dimensions to guarantee the regularity of area minimizing hypersurfaces. This proof can be extended to dimensions $\leq 10$ (equivalently, $k\leq 5$) with the recent breakthrough of Chodosh–Mantoulidis–Schulze [8, 9] on the generic regularity of miniming hypersurfaces. One may also possibly carry out the inductive descent argument with the presence of singularities in light of Schoen–Yau [28]. Note that this proof relies only on the homeomorphic type (rather than the smooth structure) of $X$.
@@ 1.5. Acknowledgment
The authors are greatly in debt to Larry Guth for stimulating conversations on the subject of this paper, and Hongbin Sun for discussions on $3$-manifolds. We also thank Misha Gromov, Bernhard Hanke and Claude LeBrun for bringing up related examples on orientable manifolds. C.L. was partially supported by an NSF grant (DMS-2202343), a Simons Junior Faculty Fellowship and a Sloan Fellowship. B.Z. was partially supported by an NSF grant (DMS-2405271) and a travel grant from the Simons Foundation.
@@ 2. The construction of $X$
We give a careful definition for the manifold $X$ when $n=4$ or $6$ as described in the introduction. Since every PL manifold with dimension $\leq 6$ has a unique smoothing [23, Theorem 2], it suffices to define $X$ as a PL manifold.
@@
The following result follows from the proof of the existence of regular neighborhoods in PL topology. See, for example, [18, Theorem 2.11(1)].
@@
By definition, if $P$ is a polygon in a PL manifold $M$, then $N\subset M$ is called a regular neighborhood of $P$ in $M$, if $N$ is a closed neighborhood of $P$ that is a codimension-0 PL submanifold of $M$ and collapses to $P$ (for the definition of collapsing, see [18, Section II.1]). Every pair of regular neighborhoods of $P$ are related to each other by a PL isotopy relative to $P\cup\partial M$ (see [25, Theorem 1.6.4(3)]).
@@
There exists a continuous map from $\hat{X}$ to $\mathbb{T}^{2k}$ whose $\mathbb{Z}_{2}$ degree is nonzero.
@@ 3. Schoen–Yau descent for non-orientable manifolds
Let $X$ be the manifold constructed in Section 2 for $k=2$ or $3$. Then by Lemma 2.10 and Corollary 3.4, the manifold $X$ does not admit any metric with PSC. On the other hand, by Lemma 2.9, the orientation double cover of $X$ admits PSC metrics. This gives the examples for $n=4$ and $6$.
@@
Recall that the construction of $X$ depends on the choices of a positive integer $L$. By Lemma 2.11, the total rank of the homology groups of $X^{4}\times(S^{1})^{n-4}$ diverges to infinity as $L\to\infty$, so we obtained infinitely many examples that are mutually non-homeomorphic. ∎
@@ 4. The enlargeability criterion and band width estimates
This is an extension of the original definition of enlargeable manifolds for the orientable case by Gromov–Lawson [11]. It is not hard to check that the notion of enlargeability for non-orientable manifolds depends only on the homotopy type of the manifold, and is preserved under taking connected sums or products. Generally, any manifold which admits a map of nonzero $\mathbb{Z}_{2}$ degree to a $\mathbb{Z}_{2}$–enlargeable manifold is itself enlargeable (cf. [19, Theorem 5.3], where one can apply the same proof for the non-orientable case by replacing the coefficients with $\mathbb{Z}_{2}$).
@@ 5. The $3$–dimensional case
In this section, we prove that examples in Theorem 1.1 do not exist in $3$-dimensions:
@@
The proof is based on Proposition 5.3 below, which extends the prime decomposition to non-orientable $3$-manifolds. Proposition 5.3 is probably known to experts in 3-manifold topology. See, for example, [16, Section 1.1, Exercies 4]. Here we give a brief proof since we were unable to find a direct reference. We will also sketch a second proof in a more geometric flavor that is closely related to [4, Section 3].
@@
We claim that $\pi_{2}(\widetilde{M_{i}^{\prime}})=0$ for all $i$. Assume the contrary, let $N$ be the kernel of the map $\pi_{2}(M_{i}^{\prime})\to\pi_{2}(\widetilde{M_{i}^{\prime}})$ given by (3). Since (3) is a surjection, we have $N\subsetneq\pi_{2}(M_{i}^{\prime})$. Note that $N$ is invariant under the $\pi_{1}(M_{i}^{\prime})$–action. So, by the generalized sphere theorem (see [17, Theorem 4.12]), there exists an embedded two-sided surface $\Sigma\subset M_{i}^{\prime}$ that is diffeomorphic to $S^{2}$ or $\mathbb{R}P^{2}$ such that the image of $\pi_{2}(\Sigma)\to\pi_{2}(M_{i}^{\prime})$ is not contained in $N$. On the other hand, by the properties of $M_{i}$, cutting open $M_{i}$ along $\Sigma$ yields at least one component $C$ such that $\widetilde{C}\cong S^{3}$, so the pre-image of $\Sigma$ in $\widetilde{M_{i}^{\prime}}$ is null-homotopic, and hence $\pi_{2}(\Sigma)\to\pi_{2}(M_{i}^{\prime})$ is contained in $N$. This gives a contradiction. As a result, we must have $\pi_{2}(\widetilde{M_{i}^{\prime}})=0$ for all $i$.
@@
We briefly indicate a more geometric proof of Proposition 5.3. This proof closely aligns with the argument of [4, Section 3]. In the place of the generalized sphere theorem [17, Theorem 4.12], we invoke the geometric result of Meeks–Yau [22]: for a closed $3$-manifold $(M,g)$, there exist finitely many conformal maps $f_{1},\cdots,f_{n}:S^{2}\to M$ with pairwise disjoint images, such that $\{f_{j}\}$ generates $\pi_{2}(M)$ as a $\pi_{1}(M)$-module, each $f_{j}$ minimizes area in its homotopy class, and each $f_{j}$ is either an embedding or a two-to-one covering to $\mathbb{R}P^{2}$. Observe that the image of each $f_{j}$ is two-sided. This is obvious if $f_{j}$ is an embedding (as $S^{2}$ is simply connected); if $f_{j}$ has an image of $\Sigma_{j}=\mathbb{R}P^{2}$, then it is also necessarily two-sided, as otherwise a tubular neighborhood of $\Sigma_{j}$ is diffeomorphic to $\mathbb{R}P^{3}\setminus\{pt\}$, and hence $\Sigma_{j}$ would be homotopically trivial and thus contradicting its minimizing property.

### v2

@@ Abstract.
We show that a closed non-orientable $3$-manifold admits a positive scalar curvature metric if and only if its orientation double cover does; however, for each $4\leq n\leq 7$, there exist infinitely many smooth non-orientable $n$-manifolds $M$ that are mutually non-homotopy equivalent, such that the orientation double cover of $M$ admits positive scalar curvature metrics, but every closed smooth manifold that is homotopy equivalent to $M$ cannot admit positive scalar curvature metrics. These examples were first introduced by Alpert-Balitskiy-Guth in the study of Urysohn widths.
To prove the nonexistence result, we extend the Schoen-Yau inductive descent approach to non-orientable manifolds. We also discuss band width estimates and the notion of enlargeability for non-orientable PSC manifolds.
@@ 1. Introduction
For each $4\leq n\leq 7$, there exist infinitely many mutually non-homotopy equivalent smooth closed non-orientable manifolds $M$ with dimension $n$ such that the following hold:
  1. (1)
    The orientation double cover of $M$ admits PSC metrics,
  2. (2)
    Every closed smooth manifold that is homotopy equivalent to $M$ cannot admit any PSC metric. In particular, $M$ does not admit any PSC metric.
Previously, there have been examples where a closed manifold $M$ does not admit any PSC metric while a suitable finite cover of $M$ does. The first such example - an exotic $\mathbb{R}P^{9}$ was discovered by Bérard Bergery in [5, Example 9.1]. LeBrun [^1] Manifolds in [21] actually depict a sign change on the Yamabe invariants after taking double cover. [21] and Hanke–Kotschick–Wehrheim [16] constructed more examples of this type for orientable $4$–manifolds (see also a subsequent work [1] and Rosenberg [25, 1.2]). Ruberman–Saveliev [26, Example 6.4] constructed a non-orientable example in dimension 4. We remark here that in all the above examples, the obstructions for PSC metrics depend on the smooth structure of $M$; except for the examples by Bergery which were in dimensions $\geq 9$, all other examples were in dimension 4, and the invariants that obstructed PSC were only defined in dimension 4. It would be interesting to further compare these approaches.
@@ 1.1. A brief description of the construction
We describe a construction for $n=4$ and $6$. More generally, taking product with a torus $\mathbb{T}^{l}$ gives more of such examples, as long as the result is of dimensions $\leq 7$. The example described here was first introduced by Alpert–Balitskiy–Guth [2, Example 11], based on a classical example by Gromov [11, $H_{1}^{\prime\prime}$].
The original construction in [2] gives a metric space with very interesting metric-geometric properties (most notably on Urysohn widths). By adding a generic perturbation, this metric space can be modified to a smooth manifold with similar properties. However, in order to understand differential geometric invariants - e.g. scalar curvature - we need to have a more precise control on the manifold. Therefore, instead of directly following the construction in [2], we will take a different approach by defining the desired manifold as a piecewise linear (PL) manifold. It is a classical result in geometric topology that every PL manifold with dimension $\leq 6$ has a unique smoothing (see [24, Theorem 2]). Working in the PL category allows us to define the desired manifold explicitly without taking any perturbations. This is crucial for the proof of the existence of PSC metrics on the orientation double cover.
@@
We will see in Section 2 that $X$ is non-orientable because the translation by $v_{2k+1}$ swaps $Z,Z^{\prime}$ and hence reverses the orientation of $X_{0}$. When $k=2,3$, define the smooth structure on $X$ to be the unique smoothing of $X$ as a PL manifold. We will show that if $M=X\times\mathbb{T}^{l}$ has dimension $\leq 7$, then every closed manifold homotopy equivalent to $M$ admits no PSC metrics, but the orientation double cover of $M$ admits PSC metrics.
@@ 1.2. Sketch of the proof
To show that every closed manifold homotopy equivalent to $M=X\times\mathbb{T}^{l}$ does not admit any PSC metric, we observe that the translations of $v_{1},\cdots,v_{n-1}$ give rises to integral cohomology classes $\alpha_{1},\cdots,\alpha_{n-1}\in H^{1}(X;\mathbb{Z})$ satisfying
@@ 1.4. Remarks on higher dimensions
In our proof of the nonexistence result for $X$, we work in low dimensions to guarantee the regularity of area minimizing hypersurfaces. This proof can be extended to dimensions $\leq 11$ ($k\leq 5$) with the recent breakthroughs [8, 9, 10] on the generic regularity of miniming hypersurfaces. One may also possibly carry out the inductive descent argument with the presence of singularities in light of Schoen–Yau [30]. Note that this proof relies only on the homotopy type of $X$.
@@ 1.5. Acknowledgment
The authors are greatly in debt to Larry Guth for stimulating conversations on the subject of this paper, and Hongbin Sun for discussions on $3$-manifolds. We also thank Misha Gromov, Bernhard Hanke, Claude LeBrun and Daniel Ruberman for bringing up related examples. C.L. was partially supported by an NSF grant (DMS-2202343), a Simons Junior Faculty Fellowship and a Sloan Fellowship. B.Z. was partially supported by an NSF grant (DMS-2405271) and a travel grant from the Simons Foundation.
@@ 2. The construction of $X$
We give a careful definition for the manifold $X$ when $n=4$ or $6$ as described in the introduction. Since every PL manifold with dimension $\leq 6$ has a unique smoothing [24, Theorem 2], it suffices to define $X$ as a PL manifold.
@@
The following result follows from the proof of the existence of regular neighborhoods in PL topology. See, for example, [19, Theorem 2.11(1)].
@@
By definition, if $P$ is a polygon in a PL manifold $M$, then $N\subset M$ is called a regular neighborhood of $P$ in $M$, if $N$ is a closed neighborhood of $P$ that is a codimension-0 PL submanifold of $M$ and collapses to $P$ (for the definition of collapsing, see [19, Section II.1]). Every pair of regular neighborhoods of $P$ are related to each other by a PL isotopy relative to $P\cup\partial M$ (see [27, Theorem 1.6.4(3)]).
@@
There exists a continuous map from $X$ to $\mathbb{T}^{2k}$ whose $\mathbb{Z}_{2}$ degree is nonzero.
@@ 3. Schoen–Yau descent for non-orientable manifolds
Let $X$ be the manifold constructed in Section 2 for $k=2$ or $3$. By comparing the top-dimensional homology groups with $\mathbb{Z}_{2}$ coefficients, every closed manifold $X^{\prime}$ homotopy equivalent to $X$ is of dimension $n=2k$. Then by Lemma 2.10 and Corollary 3.4, the manifold $X^{\prime}$ does not admit any metric with PSC. On the other hand, by Lemma 2.9, the orientation double cover of $X$ admits PSC metrics. This gives the examples for $n=4$ and $6$.
@@
Recall that the construction of $X$ depends on the choices of a positive integer $L$. By Lemma 2.11, the total rank of the homology groups of $X^{4}\times(S^{1})^{n-4}$ diverges to infinity as $L\to\infty$, so we obtained infinitely many examples that are mutually non-homotopy equivalent. ∎
@@ 4. The enlargeability criterion and band width estimates
This is an extension of the original definition of enlargeable manifolds for the orientable case by Gromov–Lawson [12]. It is not hard to check that the notion of enlargeability for non-orientable manifolds depends only on the homotopy type of the manifold, and is preserved under taking connected sums or products. Generally, any manifold which admits a map of nonzero $\mathbb{Z}_{2}$ degree to a $\mathbb{Z}_{2}$–enlargeable manifold is itself enlargeable (cf. [20, Theorem 5.3], where one can apply the same proof for the non-orientable case by replacing the coefficients with $\mathbb{Z}_{2}$).
@@ 5. The $3$–dimensional case
In this section, we prove that examples described by Theorem 1.1 do not exist in $3$-dimensions:
@@
The proof is based on Proposition 5.3 below, which extends the prime decomposition to non-orientable $3$-manifolds. Proposition 5.3 is probably known to experts in 3-manifold topology. See, for example, [17, Section 1.1, Exercies 4]. Here we give a brief proof since we were unable to find a direct reference. We will also sketch a second proof in a more geometric flavor that is closely related to [4, Section 3].
@@
We claim that $\pi_{2}(\widetilde{M_{i}^{\prime}})=0$ for all $i$. Assume the contrary, let $N$ be the kernel of the map $\pi_{2}(M_{i}^{\prime})\to\pi_{2}(\widetilde{M_{i}^{\prime}})$ given by (3). Since (3) is a surjection, we have $N\subsetneq\pi_{2}(M_{i}^{\prime})$. Note that $N$ is invariant under the $\pi_{1}(M_{i}^{\prime})$–action. So, by the generalized sphere theorem (see [18, Theorem 4.12]), there exists an embedded two-sided surface $\Sigma\subset M_{i}^{\prime}$ that is diffeomorphic to $S^{2}$ or $\mathbb{R}P^{2}$ such that the image of $\pi_{2}(\Sigma)\to\pi_{2}(M_{i}^{\prime})$ is not contained in $N$. On the other hand, by the properties of $M_{i}$, cutting open $M_{i}$ along $\Sigma$ yields at least one component $C$ such that $\widetilde{C}\cong S^{3}$, so the pre-image of $\Sigma$ in $\widetilde{M_{i}^{\prime}}$ is null-homotopic, and hence $\pi_{2}(\Sigma)\to\pi_{2}(M_{i}^{\prime})$ is contained in $N$. This gives a contradiction. As a result, we must have $\pi_{2}(\widetilde{M_{i}^{\prime}})=0$ for all $i$.
@@
We briefly indicate a more geometric proof of Proposition 5.3. This proof closely aligns with the argument of [4, Section 3]. In the place of the generalized sphere theorem [18, Theorem 4.12], we invoke the geometric result of Meeks–Yau [23]: for a closed $3$-manifold $(M,g)$, there exist finitely many conformal maps $f_{1},\cdots,f_{n}:S^{2}\to M$ with pairwise disjoint images, such that $\{f_{j}\}$ generates $\pi_{2}(M)$ as a $\pi_{1}(M)$-module, each $f_{j}$ minimizes area in its homotopy class, and each $f_{j}$ is either an embedding or a two-to-one covering to $\mathbb{R}P^{2}$. Observe that the image of each $f_{j}$ is two-sided. This is obvious if $f_{j}$ is an embedding (as $S^{2}$ is simply connected); if $f_{j}$ has an image of $\Sigma_{j}=\mathbb{R}P^{2}$, then it is also necessarily two-sided, as otherwise a tubular neighborhood of $\Sigma_{j}$ is diffeomorphic to $\mathbb{R}P^{3}\setminus\{pt\}$, and hence $\Sigma_{j}$ would be homotopically trivial and thus contradicting its minimizing property.